Starting with the equation of motion, prove that the phase space of a linear harmonic oscillator is elliptical.
1
Expert's answer
2019-05-17T11:17:00-0400

The equation of motion of a harmonic oscillator has the following form:


"\ddot{x} + \omega_0^2 x = 0"

By introducing the new variable


"y = \dot{x}"

we transform the initial 2nd order ODE into the set of the two 1st order ODEs:


"\dot{y} = - \omega_0^2 x \\
\dot{x} = y"

Excluding time, we come to the single ODE defining the phase portrait of the system:


"\frac{dy}{dx} = - \frac{\omega_0^2 x}{y}"

Performing the integration, we obtain:


"\frac{y^2}{2} + \frac{\omega_0^2 x^2}{2} = c^2"

Introducing the notations


"a^2 \equiv \frac{2 c^2}{\omega_0^2}, \quad b^2 \equiv 2 c^2,"

we obtain


"\frac{x^2}{a^2} + \frac{y^2}{b^2} = 1"

which is the canonical equation of an ellipse.